OpenStudy (anonymous):

Gerald is wrapping a gift for his grandmother . The gift is in a box that is 18 in. long 10 in. wide and 3 in. high. When he had finished How much wrapping paper will be visible civering the box?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

*covering

OpenStudy (angela210793):

U have to find the Area of the box A=2*54+2*30+2*180=528

OpenStudy (angela210793):

all i've done is added areas of 6 rectangles which are 2 by 2 equal

OpenStudy (anonymous):

how do u get the numbers?

OpenStudy (angela210793):

A1=3*10=30 A2=18*10=180 A3=18*3=54 Since we have 6 rectangles 2 of which have the A=30,two with A=54 and two with A=180 we multiply all them with 2 and then sum :)

OpenStudy (anonymous):

Thankss cuz I am like not really smmart and even I could understand that ! :)

OpenStudy (angela210793):

C'mon :P:P Ur welcome :):) w8..

OpenStudy (anonymous):

how do I metal you?

OpenStudy (angela210793):

By clicking in the blue button ''good answer'' on the right See the attachment

OpenStudy (anonymous):

Id see it?

OpenStudy (angela210793):

I drew u tht so u can understand wht rectangles i am talking abt :)

OpenStudy (anonymous):

O i seen tht thx :) but I mean the blue good answer thinngyy

OpenStudy (angela210793):

on the right of my name at the end of the gray bar...

OpenStudy (anonymous):

haa I stiill dont c it

OpenStudy (angela210793):

hmmm...take a look better

OpenStudy (anonymous):

i c it noww?

OpenStudy (angela210793):

thnx :)